OpenStudy (anonymous):

medal and fan! Shannon is managing a small zoo and has been analyzing the attendance data. Shannon has found that the number of visitors increases exponentially as the temperature increases. Shannon has also found a linear equation that models the number of people who leave the park early depending on the temperature. Describe how Shannon can combine these two functions into a new function and explain what that function would predict. Please don't copy and paste from another website. Please HELP me. Thanks.

OpenStudy (anonymous):

alrighty :)

OpenStudy (perl):

You can subtract the number of people leaving the zoo from the number of people entering the zoo to get the total amount of people in the zoo at a specific temperature. So lets say the exponential part of the attendance can be modeled by the equation \( \bf{ f(t) = a\cdot b^{t-h} +k }\). Then we have to subtract the linear part \(\bf{ g(t) = mt + c }\) because you are losing attendance. So the final result would be : \(\bf{ P(t) = (a\cdot b^{t-h}+k) - (mt + c) }\) where P stands for population in the zoo at temperature t.

OpenStudy (anonymous):

so those are like 2 different examples?

OpenStudy (perl):

yes, this would be a generic function, but you can use real data and plug it into the equation.

OpenStudy (anonymous):

ohh okay. thank you so much!!!

OpenStudy (perl):

She knows how many are coming in and how many are going out at a specific temperature, and we can assume she has the exponential and linear functions, which implies she has values where I used the constants, a,b,c,h,k,m. But t and P are variables in the equation.